Jacksonville Corp. is a U.S.‑based firm that needs $600,000. It has no business in Japan but is considering one‑year financing with Japanese yen, because the annual interest rate would be 5 percent versus 9 percent in the United States. Assume that interest rate parity exists.

a) Can Jacksonville benefit from borrowing Japanese yen and simultaneously purchasing yen one year forward to avoid exchange rate risk? Explain.

b) Assume that Jacksonville does not cover its exposure and uses the forward rate to forecast the future spot rate. Determine the expected effective financing rate. Should Jacksonville finance with Japanese yen? Explain.

c) Assume that Jacksonville does not cover its exposure and expects that the Japanese yen will appreciate by either 5 percent, 3 percent, or 2 percent, and with equal probability of each occurrence. Use this information to determine the probability distribution of the effective financing rate. Should Jacksonville finance with Japanese yen? Explain.
Answer #1

(a): No Jacksonville will not benefit from borrowing Japanese yen and simultaneously purchasing yen one year forward to avoid exchange rate risk. This is because the forward premium that it will pay will offset the interest rate differential.

Based on interest rate parity, the forward premium is about 3.8%. The effective financing rate would be: (1 + 5%)(1 + 3.8%) – 1 = about 9%

Note that (1.09/1.05) – 1 = 3.8%

(b): If it does not cover the exposure but uses the forward rate as a forecast, the expected percentage change in the Japanese yen’s value is about 3.8 percent. Thus, the expected effective financing rate is 9%. Jacksonville should therefore finance with dollars rather than Japanese yen, since the expected cost of financing with dollars is not higher.

(c ):

Possible % change in spot rate of JPY Effective financing rate of JPY if that % change occurs Probability
5% (1.05)(1.05) - 1 = 10.25% 33.30%
3% (1.05)(1.03) - 1 = 8.15% 33.30%
2% (1.05)(1.02) - 1 = 7.10% 33.30%

Given the probability, there is about a 67 percent chance that financing with Japanese yen will be less costly than financing with dollars. The choice of financing with yen or dollars in this case is dependent on Jacksonville’s degree of risk aversion.